OpenStudy (anonymous):

Just started to solve problem set 1.I need some hints regarding problem 2 part a

OpenStudy (phi):

if you are referring to the question: show the polarization identity \[ \frac{1}{4}\left( | u+v|^2 - | u-v|^2\right) = u \cdot v\] use the definition \[ | x |^2 = x \cdot x \] where in this case x is u+v or u - v expand the dot products, and simplify to get the result